3. A working substance undergoes a reversible non flow process during which P

= (-3V + 100) psia, where V changes from 10 ft3/min to 30 ft3/min. The change

of internal energy for the process is – 100 BTU/min. For a non-flow process,

determine a) work (BTU/min) b) heat transferred (BTU/min).


1
Expert's answer
2022-01-30T14:15:44-0500

"P=(-3V+100)psia"



a) Work"=\int_{V_1}^{V_2}PdV"


"=\int_{10}^{30}-3V+100dV"




"\frac{-3V^2}{2}+100V\mid_{10}^{30}"



"\frac{-3(30^2)}{2}+100(30)+\frac{3(10^2)}{2}-100(10)\\~~~~~~=800BTU/min"


b) Heat transferred"= 100-800= -700BTU\/min"


"Note: E_{int}=Q-W"